1/3(3x+9)+2=1/9(9x+18)+3

Answer by checkley77(12844)

1/3(3x+9)+2=1/9(9x+18)+3
(3x+9)/3+2=(9x+18)/9+3
(x+3)+2=(x+2)+3
x+5=x+5 proof.